Cooling Tools, Inc. is currently producing 976 of small refrigerators per month but the company’s CEO plans to increase production at a rate of 9.95 percent per month until the firm is producing 7,340 of refrigerators per month. How many months will this take?

Round the answer to two decimal places.
Answer #1

Let x be total months :

Total refrigerators to be produced = Current production of refrigerators (1 + per month rate of increment in production)x

7340 = 976 (1 + 0.0995) x

7340 = 976 (1.0995)x

7340 / 976 = (1.0995)x

7.5205 = (1.0995)x

(1.0995)21.27 = (1.0995)x

x = 21.27 months

Months will be taken for producing 7,340 of refrigerators are 21.27 months

Note:- 7.5205 = (1.0995)21.27